OpenStudy (jenniferfournierr):

will give medal if someone could pls explain steps im super confused Find f(x) and g(x) so that the function can be described as y = f(g(x)) y = eight divided by x squared' + 2

zepdrix (zepdrix):

Well hello there miss Jennifer :) Is this what the function y looks like?\[\large\rm y=\frac{8}{x^2}+2\]

OpenStudy (jenniferfournierr):

yes!

zepdrix (zepdrix):

So we need to create this y using a composition. Hmm let's be sneaky. Let's start with something like this maybe,\[\large\rm \frac{8}{x}+2\]Hmmmm

zepdrix (zepdrix):

How would we turn this expression into y? What needs to happen? We need to replace the x with something... any ideas?

OpenStudy (jenniferfournierr):

could we replace it with 2?

zepdrix (zepdrix):

We're replacing the whole x, hmm let's see what happens when we try that :)\[\large\rm \frac{8}{\color{orangered}{x}}+2\qquad\to\qquad \frac{8}{\color{orangered}{2}}+2\]Ooooo that didn't quite work out right, did it? :o

OpenStudy (jenniferfournierr):

no it didnt

OpenStudy (jenniferfournierr):

im so confused :(

zepdrix (zepdrix):

This part shouldn't be confusing :( Hmm\[\large\rm \frac{8}{~}+2\]What's missing?